OpenStudy (anonymous):

ind the particular solution y=f(x) to the given differential equation with the initial condition f(-2)=-1 dy/dx=2xe^3y

OpenStudy (anonymous):

\[\frac{ dy }{ dx}=2xe ^{3y}\]

OpenStudy (anonymous):

\[ e^{- 3y } dy= 2 x dx \]

OpenStudy (anonymous):

It is a separable equation

OpenStudy (anonymous):

\[ \int e^{- 3y } dy=\int 2 x dx \]

OpenStudy (anonymous):

Can you finish it?

OpenStudy (anonymous):